A project has an initial cost of $56,275, expected net cash inflows of $9,000 per year for 7 years, and a cost of capital of 8%. What is the project's PI? (Hint: Begin by constructing a time line.) Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Profitability index (PI)=(Present value of future cash flows)/(Initial investment)

Now, given that the net cash flows will be $9000 per year for 7 years
Present value of future cash flows=9000/(1+8%)^1+9000/(1+8%)^2+9000/(1+8%)^3+9000/(1+8%)^4+9000/(1+8%)^5+9000/(1+8%)^6+9000/(1+8%)^7

=8333.333333+7716.049383+7144.490169+6615.268675+6125.248773+5671.526642+5251.413557
=46857.33
Initial cash outflow or investment=$56275

Profitability index (PI)=(46857.33)/(56275)=0.8326 or 0.83 (Rounded to two decimal places)